OpenStudy (anonymous):

Given the system of equations, what is the solution? 2x + 3y = 2 3x - 4y = 20

OpenStudy (anonymous):

When you are solving system of equation, your goal is to eliminate a variable

OpenStudy (anonymous):

right so subtract either 2x or 3y so it can cancel out and do the same to the other aside

OpenStudy (anonymous):

Let's eliminate x, to do so we multiply first equation by 3 and second by -2 6x+9y=6 -6x+8y=-40 --------------- Now when you add the equations, the x should cancel out 17y=-34 y=2

OpenStudy (bahrom7893):

now plug y =2 back in and solve for x

OpenStudy (anonymous):

huh?

OpenStudy (bahrom7893):

y = 2,so: 2x + 3 * 2 = 2 2x + 6 = 2 2x = -4 x = -2

OpenStudy (anonymous):

so my answer is -4 and something?

OpenStudy (bahrom7893):

no it's (-2; 2)

OpenStudy (anonymous):

ok thanks had got confusedd